# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

### v2

@@ 3.3. Approximation procedure and key integral inequality
$\|f\|_{L^{\beta_{k+1}}(\Omega_{k+1})}\leq\prod_{i=1}^{k}\zeta(t_{i})^{\frac{1}{\beta_{i}}}16^{\sum_{i=1}^{k}\frac{i}{\beta_{i}}}r^{-\sum_{i=1}^{k}\frac{2}{\beta_{i}}}\|f\|_{L^{\beta_{1}}(\Omega_{k})}.$ (3.45)
@@
${\rm{vol}}(B_{r})\leq\omega_{n}r^{n}+C(n,\gamma)\|{\rm Ric}_{-}\|^{\gamma}_{L^{\gamma}}r^{2\gamma}.$ (3.46)
@@
$q\in\left[\frac{n+3}{n-1},~\frac{n+2}{(n-2)_{+}}\right).$
@@
4.1. Case 1: $n\geq 4\quad\&\quad q\in\left[\frac{n+3}{n-1},~\frac{n+2}{n-2}\right)$
4.1.1. First auxiliary function $F$ and estimation of the leading coefficients.
@@ 4.1. Case 1: $n\geq 4\quad\&\quad q\in\left[\frac{n+3}{n-1},~\frac{n+2}{n-2}\right)$
$\Delta\omega=\left(1+\frac{1}{\theta}\right)\frac{|\nabla\omega|^{2}}{\omega}+\theta\omega v^{q-1}.$ (4.1)
@@
$F=(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right).$ (4.2)
@@
$\begin{split}(v+\varepsilon)^{\theta}\Delta F=\,&2\omega^{-2}\left|\nabla^{2}\omega-\frac{\Delta\omega}{n}g\right|^{2}+2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)\\ &+2\left(\frac{1}{\theta}-\frac{\varepsilon}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\\ &+U\frac{|\nabla\omega|^{4}}{\omega^{4}}+V\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+Wv^{2(q-1)},\end{split}$ (4.3)
@@
Moreover, for $n\geq 4$ and $q\in\left[\frac{n+3}{n-1},~\frac{n+2}{n-2}\right)$, there exist $\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right)$, $d=d(n,q)>0$, $L=L(n,q)>0$ and $\widetilde{M}=\widetilde{M}(n,q)>0$ such that for any $\varepsilon>0$,
@@
$V$ $\geq V_{0}-\widetilde{M}\bar{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
$W$ $\geq W_{0}-\widetilde{M}\bar{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
where $U_{0}$, $V_{0}$, $W_{0}$ are positive constants depending only on $n$ and $q$, and $\bar{\chi}$ denotes the characteristic function.
@@
Let $n\geq 4$ and $q\in\left[\frac{n+3}{n-1},~\frac{n+2}{n-2}\right)$. Then there exist $\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right)$, $d=d(n,q)>0$, $C_{0}=C_{0}(n,q)>0$ and $M_{0}=M_{0}(n,q)>0$, such that the following holds point-wisely in $B(o,R)$:
$\Delta F\geq-2\mathrm{Ric}_{-}F-\frac{2}{\theta}|\nabla F||\nabla\ln\omega|+C_{0}(v+\varepsilon)^{\theta}F^{2}-M_{0}\varepsilon^{q-1}F.$ (4.4)
@@
$\begin{split}(v+\varepsilon)^{\theta}\Delta F\geq\,&2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)+2\left(\frac{1}{\theta}-\frac{\varepsilon}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\\ &+U_{0}\frac{|\nabla\omega|^{4}}{\omega^{4}}+V_{0}\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+W_{0}v^{2(q-1)}\\ &-\widetilde{M}v^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\bar{\chi}_{\left\{x\in B(o,R):v(x)<L\varepsilon\right\}},\end{split}$ (4.5)
@@
$2\left(\frac{1}{\theta}-1+\frac{v}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\geq-\frac{2}{\theta}(v+\varepsilon)^{\theta}|\nabla F||\nabla\ln\omega|,$ (4.6)
$\begin{split}U_{0}\frac{|\nabla\omega|^{4}}{\omega^{4}}+V_{0}\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+W_{0}v^{2(q-1)}\geq&\min\left\{U_{0},\frac{V_{0}}{2},W_{0}\right\}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)^{2}\\ \geq&C_{0}(n,q)\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)^{2}\\ =&C_{0}(n,q)(v+\varepsilon)^{2\theta}F^{2},\end{split}$ (4.7)
@@
$\begin{split}-\widetilde{M}v^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\chi_{\left\{x\in B(o,R):v(x)<L\varepsilon\right\}}\geq&-\widetilde{M}(L\varepsilon)^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\\ \geq&-M_{0}(n,q)\varepsilon^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)\\ =&-M_{0}(n,q)\varepsilon^{q-1}(v+\varepsilon)^{\theta}F,\end{split}$ (4.8)
@@
$+C_{0}(v+\varepsilon)^{\theta}F^{2}-M_{0}\varepsilon^{q-1}F.$ (4.9)
@@
Substituting the above inequality into (4.1.1), we finish the proof of Lemma 4.2. ∎
4.1.2. Integral estimate
@@
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Let $n\geq 4$, $q\in\left[\frac{n+3}{n-1},~\frac{n+2}{n-2}\right)$ and $\Omega=B(o,R)$. Define $\theta$, $d$, $C_{0}$ and $M_{0}$ as in Lemma 4.2. Then, for
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},~+\infty\right),$ (4.10)
@@
$\begin{split}&C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}+\frac{\mathbb{S}_{\chi}(M)}{2t}\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2},\end{split}$ (4.11)
@@
$\begin{split}&2\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+\frac{2}{\theta}\int_{\Omega}F^{t}|\nabla F||\nabla\ln\omega|\eta^{2}+M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}-2\int_{\Omega}F^{t}\langle\nabla F,\nabla\eta\rangle\eta\\ \geq&t\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}.\end{split}$ (4.12)
@@
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},~+\infty\right),$
@@
$\begin{split}&t\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{8}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.\end{split}$ (4.13)
@@
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Assume $v$ is a positive solution to (1.2) on the geodesic ball $B(o,R)\subset M$. Furthermore, let $n\geq 4$, $q\in\left[\frac{n+3}{n-1},~\frac{n+2}{n-2}\right)$ and $\Omega=B(o,R)$. Define $t$ and $\theta$ as in Lemma 4.3. Assume further that
@@
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}\left(\frac{1}{R^{2(t+2)}\varepsilon^{\theta(t+1)}}+\varepsilon^{(q-1)+(q-\theta-1)(t+1)}\right)V,$ (4.14)
@@
$C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}+H(t)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\leq 2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$ (4.15)
@@
$\begin{cases}0\leq\eta_{1}\leq 1,\quad\eta_{1}\equiv 1~in~B\left(o,\frac{3}{4}R\right);\\[8.53581pt] |\nabla\eta_{1}|\leq\frac{C(n)}{R}.\end{cases}$
@@
$\begin{split}\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}\leq&\frac{C^{2}(n)}{R^{2}}\frac{12(t+2)^{2}}{t}\int_{\Omega}F^{t+1}\eta^{\frac{2(t+1)}{t+2}}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\Omega}F^{t+2}\eta^{2}+C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}},\end{split}$ (4.16)
@@
$\begin{split}2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}=&2M_{0}\varepsilon^{q-1}\int_{\widetilde{\Omega}}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega\setminus\widetilde{\Omega}}F^{t+1}\eta^{2}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\widetilde{\Omega}}F^{t+2}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega\setminus\widetilde{\Omega}}\left(\frac{4M_{0}}{C_{0}}\varepsilon^{q-\theta-1}\right)^{t+1}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\Omega}F^{t+2}\eta^{2}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V,\end{split}$ (4.17)
@@
$\begin{split}H(t)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\leq C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V.\end{split}$ (4.18)
@@
Next, we choose a large $t$ such that
$t+3-\beta^{*}>0,\quad\quad\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\beta^{*}>0$ (4.19)
@@
Since $\mathrm{vol}\left(B(o,R)\right)=O(R^{\beta^{*}})$, we obtain
$H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{4}\left(\frac{1}{R^{t+3-\beta^{*}}}+\frac{1}{R^{\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\beta^{*}}}\right).$ (4.20)
@@
4.2. Case 2: $n\in\{2,3\}\quad\&\quad q\in\left[\frac{n+3}{n-1},~\frac{n+2}{(n-2)_{+}}\right)$
4.2.1. Second auxiliary function $G$ and estimation of the leading coefficients.
@@ 4.2. Case 2: $n\in\{2,3\}\quad\&\quad q\in\left[\frac{n+3}{n-1},~\frac{n+2}{(n-2)_{+}}\right)$
$\Delta\omega=\left(1+\frac{1}{\theta}\right)\frac{|\nabla\omega|^{2}}{\omega}+\theta\omega\frac{v^{q}}{v+\varepsilon}.$ (4.21)
@@
$G=(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right).$ (4.22)
@@
Moreover, for $n\in\{2,3\}$ and $q\in\left[\frac{n+3}{n-1},~\frac{n+2}{(n-2)_{+}}\right)$, there exist constants $\theta=\theta(n,q)\in\left(0,\,q-1\right)$ if $n=2$ or $\theta=\theta(n,q)\in\left(0,\,\min\left\{2,q-1\right\}\right)$ if $n=3$, $d=d(n,q)>0$, $L=L(n,q)>0$ and $\widetilde{M}=\widetilde{M}(n,q)>0$ such that for any $\varepsilon>0$,

### v3

@@ 3.C. Approximation procedure and key integral inequality
(3.45) $\|f\|_{L^{\beta_{k+1}}(\Omega_{k+1})}\leq\prod_{i=1}^{k}\zeta(t_{i})^{\frac{1}{\beta_{i}}}16^{\sum_{i=1}^{k}\frac{i}{\beta_{i}}}r^{-\sum_{i=1}^{k}\frac{2}{\beta_{i}}}\|f\|_{L^{\beta_{1}}(\Omega_{k})}.$
@@
(3.46) ${\rm{Vol}}(B_{r})\leq\omega_{n}r^{n}+C(n,\gamma)\|\mathrm{Ric}_{-}\|^{\gamma}_{L^{\gamma}}r^{2\gamma}.$
@@
$q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{(n-2)_{+}}\right).$
@@
4.A. Case 1: $n\geq 4\quad\&\quad q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$
4.A.1. First auxiliary function $F$ and estimation of the leading coefficients.
@@ 4.A. Case 1: $n\geq 4\quad\&\quad q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$
(4.1) $\Delta\omega=\left(1+\frac{1}{\theta}\right)\frac{|\nabla\omega|^{2}}{\omega}+\theta\omega v^{q-1}.$
@@
(4.2) $F=(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right).$
@@
(4.3) $\begin{split}(v+\varepsilon)^{\theta}\Delta F=\,&2\omega^{-2}\left|\nabla^{2}\omega-\frac{\Delta\omega}{n}g\right|^{2}+2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)\\ &+2\left(\frac{1}{\theta}-\frac{\varepsilon}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\\ &+U\frac{|\nabla\omega|^{4}}{\omega^{4}}+V\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+Wv^{2(q-1)},\end{split}$
@@
Moreover, for $n\geq 4$ and $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$, there exist $\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right)$, $d=d(n,q)>0$, $L=L(n,q)>0$ and $\widetilde{M}=\widetilde{M}(n,q)>0$ such that for any $\varepsilon>0$,
@@
$V$ $\geq V_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
$W$ $\geq W_{0}-\widetilde{M}\overline{\chi}_{\left\{x\in B(o,R):\,v(x)<L\varepsilon\right\}},$
where $U_{0}$, $V_{0}$, $W_{0}$ are positive constants depending only on $n$ and $q$, and $\overline{\chi}$ denotes the characteristic function.
@@
Let $n\geq 4$ and $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$. Then there exist $\theta=\theta(n,q)\in\left(0,\frac{2}{n-2}\right)$, $d=d(n,q)>0$, $C_{0}=C_{0}(n,q)>0$ and $M_{0}=M_{0}(n,q)>0$, such that the following holds point-wisely in $B(o,R)$:
(4.4) $\Delta F\geq-2\mathrm{Ric}_{-}F-\frac{2}{\theta}|\nabla F||\nabla\ln\omega|+C_{0}(v+\varepsilon)^{\theta}F^{2}-M_{0}\varepsilon^{q-1}F.$
@@
(4.5) $\begin{split}(v+\varepsilon)^{\theta}\Delta F\geq\,&2\omega^{-2}\mathrm{Ric}(\nabla\omega,\nabla\omega)+2\left(\frac{1}{\theta}-\frac{\varepsilon}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\\ &+U_{0}\frac{|\nabla\omega|^{4}}{\omega^{4}}+V_{0}\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+W_{0}v^{2(q-1)}\\ &-\widetilde{M}v^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\overline{\chi}_{\left\{x\in B(o,R):v(x)<L\varepsilon\right\}},\end{split}$
@@
(4.6) $2\left(\frac{1}{\theta}-1+\frac{v}{v+\varepsilon}\right)(v+\varepsilon)^{\theta}\langle\nabla F,\nabla\ln\omega\rangle\geq-\frac{2}{\theta}(v+\varepsilon)^{\theta}|\nabla F||\nabla\ln\omega|,$
(4.7) $\begin{split}U_{0}\frac{|\nabla\omega|^{4}}{\omega^{4}}+V_{0}\frac{|\nabla\omega|^{2}}{\omega^{2}}v^{q-1}+W_{0}v^{2(q-1)}\geq&\min\left\{U_{0},\frac{V_{0}}{2},W_{0}\right\}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)^{2}\\ \geq&C_{0}(n,q)\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)^{2}\\ =&C_{0}(n,q)(v+\varepsilon)^{2\theta}F^{2},\end{split}$
@@
(4.8) $\begin{split}-\widetilde{M}v^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\chi_{\left\{x\in B(o,R):v(x)<L\varepsilon\right\}}\geq&-\widetilde{M}(L\varepsilon)^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+v^{q-1}\right)\\ \geq&-M_{0}(n,q)\varepsilon^{q-1}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right)\\ =&-M_{0}(n,q)\varepsilon^{q-1}(v+\varepsilon)^{\theta}F,\end{split}$
@@
(4.9) $+C_{0}(v+\varepsilon)^{\theta}F^{2}-M_{0}\varepsilon^{q-1}F.$
@@
Substituting the above inequality into (4.A.1), we finish the proof of Lemma 4.2. ∎
4.A.2. Integral estimate
@@
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Let $n\geq 4$, $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$ and $\Omega=B(o,R)$. Define $\theta$, $d$, $C_{0}$ and $M_{0}$ as in Lemma 4.2. Then, for
(4.10) $t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},\penalty 10000\ +\infty\right),$
@@
(4.11) $\begin{split}&C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}+\frac{\mathbb{S}_{\chi}(M)}{2t}\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2},\end{split}$
@@
(4.12) $\begin{split}&2\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+\frac{2}{\theta}\int_{\Omega}F^{t}|\nabla F||\nabla\ln\omega|\eta^{2}+M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}-2\int_{\Omega}F^{t}\langle\nabla F,\nabla\eta\rangle\eta\\ \geq&t\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}.\end{split}$
@@
$t\in\left(\max\left\{\frac{8}{C_{0}\theta^{2}},1\right\},\penalty 10000\ +\infty\right),$
@@
(4.13) $\begin{split}&t\int_{\Omega}F^{t-1}|\nabla F|^{2}\eta^{2}+C_{0}\int_{\Omega}(v+\varepsilon)^{\theta}F^{t+2}\eta^{2}\\ \leq&4\int_{\Omega}\mathrm{Ric}_{-}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{8}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.\end{split}$
@@
Let $(M,g)$ be a complete manifold on which the $\chi$-type Sobolev inequality holds. Assume $v$ is a positive solution to (1.2) on the geodesic ball $B(o,R)\subset M$. Furthermore, let $n\geq 4$, $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{n-2}\right)$ and $\Omega=B(o,R)$. Define $t$ and $\theta$ as in Lemma 4.3. Assume further that
@@
(4.14) $H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{3}\left(\frac{1}{R^{2(t+2)}\varepsilon^{\theta(t+1)}}+\varepsilon^{(q-1)+(q-\theta-1)(t+1)}\right)V,$
@@
(4.15) $C_{0}\varepsilon^{\theta}\int_{\Omega}F^{t+2}\eta^{2}+H(t)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\leq 2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}+\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}.$
@@
$\begin{cases}0\leq\eta_{1}\leq 1,\quad\eta_{1}\equiv 1\penalty 10000\ in\penalty 10000\ B\left(o,\frac{3}{4}R\right);\\[8.53581pt] |\nabla\eta_{1}|\leq\frac{C(n)}{R}.\end{cases}$
@@
(4.16) $\begin{split}\frac{12}{t}\int_{\Omega}F^{t+1}|\nabla\eta|^{2}\leq&\frac{C^{2}(n)}{R^{2}}\frac{12(t+2)^{2}}{t}\int_{\Omega}F^{t+1}\eta^{\frac{2(t+1)}{t+2}}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\Omega}F^{t+2}\eta^{2}+C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}},\end{split}$
@@
(4.17) $\begin{split}2M_{0}\varepsilon^{q-1}\int_{\Omega}F^{t+1}\eta^{2}=&2M_{0}\varepsilon^{q-1}\int_{\widetilde{\Omega}}F^{t+1}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega\setminus\widetilde{\Omega}}F^{t+1}\eta^{2}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\widetilde{\Omega}}F^{t+2}\eta^{2}+2M_{0}\varepsilon^{q-1}\int_{\Omega\setminus\widetilde{\Omega}}\left(\frac{4M_{0}}{C_{0}}\varepsilon^{q-\theta-1}\right)^{t+1}\\ \leq&\frac{C_{0}\varepsilon^{\theta}}{2}\int_{\Omega}F^{t+2}\eta^{2}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V,\end{split}$
@@
(4.18) $\begin{split}H(t)\left\|F^{t+1}\eta^{2}\right\|_{L^{\chi}(\Omega)}\leq C_{1}(n,q,t)\varepsilon^{-\theta(t+1)}\frac{V}{R^{2(t+2)}}+C_{2}(n,q,t)\varepsilon^{(q-1)+(q-\theta-1)(t+1)}V.\end{split}$
@@
Next, we choose a large $t$ such that, as $\chi\in(1,\frac{n}{n-2})$,
(4.19) $t+3-\frac{2\chi}{\chi-1}>0\quad\mbox{and}\quad\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\frac{2\chi}{\chi-1}>0,$
or as $\chi=\frac{n}{n-2}$,
(4.20) $t+3-\beta^{*}>0\quad\mbox{and}\quad\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\beta^{*}>0,$
@@
Since
$\mathrm{vol}\left(B(o,R)\right)=O(R^{\frac{2\chi}{\chi-1}})\quad\mbox{or}\quad\mathrm{vol}\left(B(o,R)\right)=O(R^{\beta^{*}}),$
by setting $\beta^{*}_{\chi}=\frac{2\chi}{\chi-1}$ as $1<\chi<\frac{n}{n-2}$ and $\beta^{*}_{\chi}=\beta^{*}$ as $\chi=\frac{n}{n-2}$, we obtain
(4.21) $H(t)\left\{\int_{B\left(o,\frac{3}{4}R\right)}\left[(v+1)^{-\theta}v^{q-1}\right]^{(t+1)\chi}\right\}^{\frac{1}{\chi}}\leq C_{4}\left(\frac{1}{R^{t+3-\beta^{*}_{\chi}}}+\frac{1}{R^{\frac{1}{\theta}\left[(q-1)+(q-\theta-1)(t+1)\right]-\beta^{*}_{\chi}}}\right).$
@@
4.B. Case 2: $n\in\{2,3\}\quad\&\quad q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{(n-2)_{+}}\right)$
4.B.1. Second auxiliary function $G$ and estimation of the leading coefficients.
@@
(4.22) $\Delta\omega=\left(1+\frac{1}{\theta}\right)\frac{|\nabla\omega|^{2}}{\omega}+\theta\omega\frac{v^{q}}{v+\varepsilon}.$
@@
(4.23) $G=(v+\varepsilon)^{-\theta}\left(\frac{|\nabla\omega|^{2}}{\omega^{2}}+dv^{q-1}\right).$
@@
Moreover, for $n\in\{2,3\}$ and $q\in\left[\frac{n+3}{n-1},\penalty 10000\ \frac{n+2}{(n-2)_{+}}\right)$, there exist constants $\theta=\theta(n,q)\in\left(0,\,q-1\right)$ if $n=2$ or $\theta=\theta(n,q)\in\left(0,\,\min\left\{2,q-1\right\}\right)$ if $n=3$, $d=d(n,q)>0$, $L=L(n,q)>0$ and $\widetilde{M}=\widetilde{M}(n,q)>0$ such that for any $\varepsilon>0$,